# ATRNet-STAR: A Large Dataset and Benchmark Towards Remote Sensing Object Recognition in the Wild

### v3

@@ 3.4 Statistical Analysis
Class distribution. ATRNet-STAR includes 4 class, 21 subclasses and 40 types of vehicles with balanced samples, as shown in Fig. 5. Relative to extant SAR target datasets, our dataset’s targets’ enhanced diversity and intricacy introduce novel formidable challenges to the SAR target fine-grained recognition research. Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
@@
Object size. The length, width, and height of our 40 objects are listed in Fig. 7, and we can see that they have various structures, different sizes, and similar size ratios in Fig. 8. Compared to another civilian vehicle dataset (Gotcha), our dataset has a wider range of target classes and size distribution.
@@
Reference target. Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering issues of a standard geometry.
@@
Non-centered location. Most SAR target datasets place the target in the image center or contain only the target region, but the QinetiQ dataset uses a non-centered setting. Since remote sensing differs from the human eye view, which customarily centers the object of interest, the overhead view requires searching targets with more interference. As shown in Fig. 9, we randomly added position offsets to increase the recognition difficulty with detecting non-centered target locations.
@@
Occlusion in scenes. Whereas previous vehicle datasets prioritize target characteristics in ideal environments, we systematically acquire samples across diverse scenes. Fig. 10 shows their different influences with occlusions and layovers in different scene regions. The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
@@
Imaging angle. Different imaging angles affect the target scattering characteristics, and the imaging geometry relationship can change the interference in the scene. Our dataset has various imaging angles with balance distribution in different depression angles, as shown in Fig. 11. However, azimuth angle sampling density exhibits scene-dependent due to stripmap mode constraints and operational cost limitations.
@@
Band and polarization. Table II shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings. However, Fig. 12 demonstrates that this setting is not the only one to consider, as band and polarization have a noticeable effect on target scattering. We provide two bands and quad polarization to support discussing these sensor parameters.
@@
Complex data. Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format.
@@
Number of image bits. RGB images are stored in 8-bit format, whereas SAR images have a larger range of digital data values. Therefore, we provide different bit formats to investigate quantification for enhancing weaker target scattering points.
@@
Cross-platform problem. We collect a small amount of satellite data synchronously under the same scenes and targets to investigate the difficulty of recognition across platforms.
@@
Quality control. To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors perform the data annotation about four months. The entire project cost nearly two years.
@@ 3.5 Dataset value
Based on the diversity of target categories, scenes, imaging conditions, and data format with detailed annotation, we recommend that this dataset be used for SAR target recognition studies in the following areas.
Robustness recognition. Our dataset encompasses diverse acquisition scenarios, with each image accompanied by exhaustive acquisition metadata. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., , scene and sensor sensors) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [100].
Few-shot learning. Besides acquisition condition variation, the inherent challenge of limited training samples due to high SAR collection costs also presents critical research opportunities. The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [101, 102, 103, 104] settings in existing few-shot learning.
Transfer learning. Our dataset contains a large diversity of samples that can be used to study the transfer problem of the pre-training model with different SAR target datasets. In addition, the large number of samples can further increase the volume of SAR target samples to advance the study of self-supervised learning [105, 29] and foundation models [106, 30].
Incremental learning. The diversity of our dataset supports investigations into domain incremental learning and class incremental learning [107]. The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes. These incremental capacities are essential for SAR ATR in open environments.
Physical deep learning. SAR images have unique properties, such as complex phase and polarization. Our dataset provides multi-format data to facilitate recognition studies leveraging these intrinsic attributes instead of only relying on quantized SAR magnitude images. In addition, detailed metadata can also provide more gains for SAR ATR.
Generative Models. Beyond recognition tasks, our dataset enables controllable generation of target samples under diverse imaging conditions [2], as well as estimation of target parameters across varying acquisition scenarios.
@@ 4 Benchmarks
In conclusion, we evaluate our benchmarks under 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images. These experiments provide valuable insights and substantial potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes. Our dataset is sufficient to provide a new, challenging benchmark for advancing SAR target recognition technology. We also welcome researchers to contact us to construct richer experimental benchmarks from original data.
@@ 4.1 Experimental settings
SOC and EOC settings. SOC settings are those where the training and test sets have similar imaging conditions, and we have SOC-40 and SOC-50 settings. SOC-40 is created from a random data sample with a 7:3 training-to-test ratio. SOC-50 randomly selects data with a similar amount of MSTAR ten classes, and we combine our dataset with MSTAR. EOC settings are those where there is a significant domain shift between the training set and the test set, such as variation in imaging conditions and target state. For EOC-Scene, we use simple scenes (sandstone and bare soil) as the training set and complex scenes (urban, factory, and woodland) with occlusion as the test set. EOC-Depression is the training set with a 15∘ depression angle and the test set with 30∘, 45∘ and 60∘. EOC-Azimuth means that training has limited target azimuth angles within 0∘60∘, and the test has other angles. Imaging angle and geometry variation can change target signatures and background clutter. EOC-band and EOC-polarization are designed to test the effect of different bands and change from HH-polarization to other polarizations on recognition, respectively. We do not consider some experimental settings that added simulation perturbations, such as random/Gaussian noise and simulation occlusions.
Compared methods. We select classical or recent open-source work in computer vision and SAR target recognition, in addition to recently proposed methods from our laboratory. For classification, we have chosen 6 computer vision methods (VGG16 [108], ResNet18 [109], ResNet34 [109], ConvNeXt [64], ViT [67], and HiViT [110]) and 4 SAR ATR methods (HDANet [53], SARATR-X [30, 29], MS-CVNets [111], and LDSF [90]). MS-CVNets and LDSF are complex images methods others are magnitude methods. VGG16, ResNet18, ResNet34, ConvNeXt, ViT, HiViT, and SARATR-X all use the pre-training weights. For detection, we have chosen 4 computer vision methods (Faster RCNN [112], CenterNet [113], YOLOv8 [114], and YOLOv10 [115]) and 2 SAR ATR methods (DiffDet4SAR [12], and SARATR-X [30, 29]). These methods are based on magnitude images. Because of the long-term dataset construction process, we will keep updating the dataset homepage with new experimental settings and method results.
Transfer learning. To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table V, and the weights are used for other SAR target classification datasets as initialization for fine-tuning. Fine-tuning parameter settings follow our previous paper [29], and the baseline is the weight of ImageNet supervised pre-training for these models.
@@ 4.2 Results and analyses
| SOC-40 | ||||||||||
| SOC-50 | ||||||||||
| EOC-Scene | ||||||||||
| – (city) | ||||||||||
| – (factory) | ||||||||||
| – (woodland) | ||||||||||
| EOC-Depression | ||||||||||
| – (30) | ||||||||||
| – (45) | ||||||||||
| – (60) | ||||||||||
| EOC-Azimuth | 15.7 | 16.4 | 14.9 | 16.5 | 21.1 | 22.8 | 26.4 | 19.0 | 22.5 | |
| – (60) | 20.7 | 19.4 | 18.0 | 22.1 | 26.2 | 27.2 | 28.9 | 26.8 | 31.0 | |
@@
Classification. From different methods in Table V, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40. Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene. The model recognition performance and robustness of the latest 2020s CNN architecture ConvNeXt are superior to the previous CNN models. ConvNeXt is more robust under EOC-Deprssion than other methods, and ViT is insensitive to azimuthal angle change. Our previously proposed method, HDANet, improves the attention mechanism and uses a segmentation task so it can be used in different scenarios. In addition, our proposed LDSF introduces a scattering graph structure that effectively improves the discrimination of targets and clutter. SARATR-X is a pre-trained self-supervised learning model on SAR with the HiViT architecture, and its performance is significantly improved compared to HiViT. Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
@@
Detection. As illustrated in Table VI, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions. In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions. However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions. In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios. This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency. The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
@@
Transfer learning. Fig 14 shows how our dataset can help different models benefit from feature learning and be used for target recognition in other datasets. We use Table’s SOC-40 trained model weights as initializations on other target datasets with 5-shot sample experiments [29]. The fine-tuning results show that similar ground target-type datasets, such as vehicles and airplanes, would greatly benefit from training on our data. However, the transfer effect is not obvious for the ship target class dataset in the sea scene. In recent years, the number of emerging satellite ship datasets has been much larger than that of other ground targets. This data can greatly compensate for the lack of high-resolution ground target samples in the existing SAR target datasets and advance the construction of an SAR ATR foundation model.
@@ 4.3 Discussion
Researchers have endeavored to realize a stable and efficient SAR ATR recognition system, but previous vehicle benchmark datasets have not provided sufficiently challenging conditions. Our results show that representative methods encounter significant challenges, such as imaging angle and scene variation, as well as few-shot settings. Robust feature extraction under different varying imaging conditions remains a critical research problem. It is also crucial to efficiently utilize a small number of labeled samples and mine complex and polarization information. Although the pre-trained foundation model SARATR-X is effective in most cases compared to other methods, no single method can effectively handle diverse perturbation conditions, highlighting the complexity of SAR ATR. In a data-driven approach dominant in the 2020s, our transfer learning results show that a large-scale dataset will help to extract stabilizing features and enhance performance across multiple datasets and tasks. Based on the results, we discuss several potential directions for the future development of ATR systems.
Robustness under different conditions. As shown in Table V and VI, the variations in imaging conditions present a significant challenge for recognition. This variability in the domain distribution requires the extraction of stable invariant target features under different conditions.
Efficient utilization of samples. Given the challenges associated with acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible light dataset. Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 13.
Incorporation with multi-dimensional information. Table V show that complex information can help distinguish between targets and clutter under scene variations. Auxiliary information such as complex and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity.
Fast incremental learning capability. Because of the open real-world ATR systems face constantly changing imaging conditions and target types. They need to effectively recognize anomalies and update their capabilities in response to variations. This dataset provides a rich set of target types and imaging conditions that can be used to explore this possibility.

### v4

@@ 3.4 Statistical Analysis
Class distribution - ATRNet-STAR dataset includes 4 classes, 21 subclasses and 40 types of vehicles with balanced samples, as shown in Fig. 5. Relative to other SAR target datasets, our dataset’s enhanced diversity introduces novel formidable challenges to the SAR target fine-grained recognition research. Moreover, adequate samples guarantee the availability of a substantial array of target classes across experimental settings, facilitating more diverse investigations.
@@
Object size - The length, width, and height of our 40 objects are listed in Fig. 7, and we can see that they have various structures, different sizes, and similar size ratios. Compared to another civilian vehicle dataset (Gotcha), our dataset has a wider range of target classes and size distribution in Fig. 8.
@@
Reference target - Besides deploying corner reflectors for resolution measurements, we produce and place a reference target of multiple geometries as the “SLICY” target in the MSTAR dataset for scattering research.
@@
Non-centered location - Most SAR target datasets place the target in the image center or contain only the target region. Since remote sensing differs from the human eye view that customarily centers the object of interest, the overhead view requires searching targets with more interference. We randomly added position offsets to increase the recognition difficulty with detecting non-centered target locations in Fig. 9.
@@
Occlusion in scenes - Whereas previous vehicle datasets collect target samples in ideal environments, we systematically acquire samples across diverse scenes. Fig. 10 shows their different influences with occlusions and layovers in different scene regions. The shadow of buildings and roadside trees in the factory may obscure targets and reduce target scattering, and the nearby trees in woodland have more obvious occlusion and increase non-target scattering.
@@
Imaging angle - Different imaging angles affect the target scattering characteristics, and the imaging geometry relationship can change the interference in the scene. Our dataset has various imaging angles with balance distribution in different depression angles across scenes, as shown in Fig. 11. However, azimuth angle sampling density exhibits scene-dependent due to stripmap mode constraints and operational cost limitations.
@@
Band and polarization - Table II shows most vehicle datasets with the MSTAR dataset’s X-band and HH polarization settings. However, Fig. 12 demonstrates that this setting is not the only one to consider, as band and polarization have a noticeable effect on target scattering. We provide two bands and quad polarization to support discussing these sensor parameters.
@@
Complex data - Feature extraction based on complex data is a hot research concern, so we provide complex data in slant distance coordinates stored in .mat format.
@@
Number of image bits - RGB images are stored in 8-bit format, whereas SAR images have a larger range of digital data values. Therefore, we provide different bit formats to investigate quantification for enhancing weaker target scattering points.
@@
Cross-platform problem - We collect a small amount of satellite data synchronously under the same scenes and targets to investigate the difficulty across platforms for future versions.
@@
Quality control - To ensure high-quality data, we considered challenges such as target and site hire, airspace applications, weather conditions, target placing, and data checking. Eight people completed the data acquisition over six months. In addition, 14 labelers and two inspectors performed the data annotation for about four months. The entire project cost nearly two years.
@@ 3.5 Dataset value
Based on the diversity of target classes, scenes, imaging conditions, and data format with detailed annotation, we recommend that this dataset be used for SAR studies in the following areas.
Robustness recognition - Our dataset encompasses diverse acquisition scenarios, with each image accompanied by exhaustive acquisition metadata. Therefore, we can discuss in detail the robustness of the SAR recognition algorithms when the acquisition conditions (e.g., , scene and sensor sensors) vary between the training and test sets. If we consider data from different acquisition conditions as domains, the training set can include multiple domains to learn domain invariant features, i.e., the domain generalization [101].
Few-shot learning - Besides acquisition condition variation, the inherent challenge of limited training samples due to high SAR collection costs also presents critical research opportunities. The few-shot setting can be combined with robust recognition, and our dataset’s large number of target types can enrich the task construction of meta-learning [102, 103, 104, 105] settings in existing few-shot learning.
Transfer learning - Our dataset contains a large diversity of samples that can be used to study the transfer problem of the pre-training model with different SAR target datasets. In addition, the large number of samples can further increase the volume of SAR target samples to advance the study of self-supervised learning [106, 30] and foundation models [107, 31].
Incremental learning - The diversity of our dataset supports investigations into domain incremental learning and class incremental learning [108]. The domain incremental learning can improve the algorithm’s robustness with a dynamic process, and class incremental learning can progressively increase the ability to recognize or reject new classes. These incremental capacities are essential for SAR ATR in open environments.
Physical deep learning - SAR images have unique properties, such as complex phase and polarization. Our dataset provides multi-format data to facilitate recognition studies leveraging these intrinsic attributes instead of only relying on quantized SAR magnitude images. In addition, detailed metadata can also provide more gains for SAR ATR.
Generative Models - Beyond recognition tasks, our dataset enables controllable generation of target samples under diverse imaging conditions [2], as well as estimation of target parameters across varying acquisition scenarios.
@@ 4.1 Experimental settings
SOC and EOC settings - SOC settings are those where the training and test sets have similar imaging conditions, and we have SOC-40 and SOC-50 settings. SOC-40 is created from a random data sample with a 7:3 training-to-test ratio. SOC-50 randomly selects data with a similar amount of MSTAR ten classes, and we combine our dataset with MSTAR. EOC settings are those where there is a significant domain shift between the training set and the test set, such as variation in imaging conditions and target state. For EOC-Scene, we use simple scenes (sandstone and bare soil) as the training set and complex scenes (urban, factory, and woodland) with occlusion as the test set. EOC-Depression is the training set with a 15∘ depression angle and the test set with 30∘, 45∘ and 60∘. EOC-Azimuth means that training has limited target azimuth angles within 0∘60∘, and the test has other angles. Imaging angle and geometry variation can change target signatures and background clutter. EOC-band and EOC-polarization are designed to test the effect of different bands and change from HH-polarization to other polarizations on recognition, respectively. We do not consider experimental settings that added simulation perturbations, such as random/Gaussian noise and simulation occlusions.
Compared methods - We select classical or recent open-source work in computer vision and SAR target recognition, in addition to recently proposed methods from our laboratory. For classification, we have chosen six computer vision methods (VGG16 [109], ResNet18 [110], ResNet34 [110], ConvNeXt [65], ViT [68], and HiViT [111]) and four SAR ATR methods (HDANet [54], SARATR-X [31, 30], MS-CVNets [112], and LDSF [91]). MS-CVNets and LDSF are complex images methods others are magnitude methods. VGG16, ResNet18, ResNet34, ConvNeXt, ViT, HiViT, and SARATR-X all use the pre-training weights. For detection, we have chosen four computer vision methods (Faster RCNN [113], CenterNet [114], YOLOv8 [115], and YOLOv10 [116]) and two SAR ATR methods (DiffDet4SAR [12], and SARATR-X [31, 30]). These methods are based on magnitude images. Because of the long-term dataset and benchmark construction process, we will keep updating the homepage with new experimental settings and method results.
Transfer learning - To show that this dataset can improve the model’s performance for SAR target feature extraction, we perform training on SOC-40 according to Table V, and the weights are used for other SAR target classification datasets as initialization for fine-tuning. Fine-tuning parameter settings follow our previous paper [30], and the baseline is the weight of ImageNet supervised pre-training for these models.
@@ 4.2 Results and analyses
| SOC-40 | 96.4 | |||||||||
| SOC-50 | 85.2 | |||||||||
| EOC-Scene | 33.6 | |||||||||
| – (city) | 33.7 | |||||||||
| – (factory) | 34.3 | |||||||||
| – (woodland) | 27.6 | |||||||||
| EOC-Depression | 43.1 | |||||||||
| – (30) | 63.2 | |||||||||
| – (45) | 45.1 | |||||||||
| – (60) | 21.3 | |||||||||
| EOC-Azimuth | 15.7 | 16.4 | 14.9 | 16.5 | 21.1 | 29.0 | 22.8 | 26.4 | 19.0 | 22.5 |
| – (60) | 20.7 | 19.4 | 18.0 | 22.1 | 26.2 | 34.4 | 27.2 | 28.9 | 26.8 | 31.0 |
@@
Classification - From different methods in Table V, we note that deep learning methods can achieve high performance in SAR fine-grained classification, such as the recent methods ConvNeXt and SARATR-X in SOC-40. Methods based on complex images, such as MS-CVNet and LDSF, better distinguish between targets and clutter in EOC-Scene. The model recognition performance and robustness of the latest 2020s CNN architecture ConvNeXt are superior to the previous CNN models. ConvNeXt is more robust under EOC-Deprssion than other methods, and ViT is insensitive to azimuthal angle change. Our previously proposed method, HDANet, improves the attention mechanism and uses a segmentation task so it can be used in different scenarios. In addition, our proposed LDSF introduces a scattering graph structure that effectively improves the discrimination of targets and clutter. SARATR-X is a pre-trained self-supervised learning model on SAR with the HiViT architecture, and its performance is significantly improved compared to HiViT. Besides, the Transformer models are trained with some low-layer weights frozen, as we found that the full update tends to lead to overfitting.
@@
Detection - As illustrated in Table VI, we find that the two-stage approach (Faster RCNN, DiffDet4SAR, and SARATR-X) has better robustness than the single-stage approach (CenterNet, YOLOv8, and YOLOv10) under most EOC conditions. In particular, CenterNet models the target center using a Gaussian kernel, demonstrating superior detection capabilities under EOC-Depression and EOC-Polarization conditions. However, YOLO series methods, which rely on fixed grid partitioning and coordinate regression, exhibit limitations in detection performance under EOC conditions. In contrast, our proposed DiffDet4SAR, a diffusion model-based detection framework, consistently outperforms other algorithms in EOC-SCENE and EOC-Azimuth scenarios. This advantage stems from the robust denoising box structure, which enhances target location accuracy, and the use of central difference convolution to amplify target saliency. The foundation model SARATR-X, based on self-supervised learning with SAR images, performs well for fine-grained recognition and most of the EOC conditions, but the performance for scene and angle changes needs to be improved.
@@
Transfer learning - Fig 14 shows how our dataset can help different models benefit from feature learning and be used for target recognition in other datasets. We use Table’s SOC-40 trained model weights as initializations on other target datasets with 5-shot sample experiments [30]. The fine-tuning results show that similar ground target-type datasets, such as vehicles and airplanes, would greatly benefit from training on our data. However, the transfer effect is not obvious for the ship target class dataset in the sea scene. In recent years, the number of emerging satellite ship datasets has been much larger than that of other ground targets. This dataset can greatly compensate for the lack of high-resolution ground target samples in the existing SAR target datasets and advance the construction of SAR ATR foundation models.
In conclusion, we evaluate our benchmarks based on 7 experimental settings and 15 recognition methods with classification and detection tasks based on magnitude images and classification tasks based on complex images. These experiments provide valuable insights and substantial potential into recognizing SAR target characteristics under different imaging conditions and fine-grained classes. Our dataset is sufficient to provide a new, challenging benchmark for advancing SAR target recognition technology. We also welcome researchers to contact us to construct richer experimental benchmarks from original data.
@@ 4.3 Discussion
Researchers have endeavored to realize a stable and efficient SAR ATR system, but previous vehicle benchmark datasets have not provided sufficiently challenging conditions. Our results show that representative methods encounter significant challenges, such as imaging angle and scene variation, as well as few-shot settings. Robust feature extraction under different varying imaging conditions remains a critical research problem. It is also crucial to efficiently utilize a small number of labeled samples and mine complex and polarization information. Although the pre-trained foundation model SARATR-X is effective in most cases compared to other methods, no single method can effectively handle diverse perturbation conditions, highlighting the complexity of SAR ATR. In a data-driven approach dominant in the 2020s, our transfer learning results show that a large-scale dataset will help to extract stabilizing features and enhance performance across multiple datasets and tasks. Based on the results, we discuss several potential directions for the future development of ATR systems.
Robustness under different conditions - Imaging condition variations present a significant challenge for recognition in Table V. This variability in the domain distribution requires the extraction of stable invariant target features across conditions.
Efficient utilization of samples - Given the challenges associated with acquiring and labeling SAR images, it is unlikely that the sample size will reach the same size as the visible-light dataset. Therefore, it is important to improve learning efficiency by utilizing a small number of samples in Fig 13.
Incorporation with multi-dimensional information - Table V show that complex information can help distinguish between targets and clutter under scene variations. Auxiliary information such as complex and angle can be introduced in SAR ATR, but increasing the dimensionality of the information may increase the instability and complexity.
Fast incremental learning capability - Because of the open real world, ATR systems face constantly changing imaging conditions and target types. They need to effectively recognize anomalies and update their capabilities in response to variations. This dataset provides a rich set of target types and imaging conditions that can be used to explore this possibility.
Explainable, interpretable, and trustworthy recognition model - Although data-driven models have become a mainstream approach for SAR ATR, the requirement for black-box models’ interpretability during the recognition process makes it difficult to meet the trust needed by actual systems. Therefore, in addition to improving recognition performance, a transparent and trustworthy recognition model is also a problem that needs to be studied [118, 42, 80].